In segment AC, the midpoint is B. If segments AC = 5x-9 and AB = 2x, what is the measure of segment AC.
Draw the line clearly labeling all points and adding the given measures.
(Someone help!)

Answer by ikleyn(52790)

.


    -------|-----------|-----------|--------

           A           B           C


Since point B is the midpoint of segment AC,

the length of segment AC is two times the length of segment AB :


    5x - 9 = 2*(2x).


Simplify and solve for x


    5x - 9 = 4x

    5x - 4x = 9

     x      = 9


The length of segment AC is  5x-9 = 5*9 - 9 = 45 - 9 = 36.    ANSWER